# EFSA: Episodic Few-Shot Adaptation for Text-to-Image Retrieval

### v1

@@ 7 Conclusion
In this paper, we argue that the text-to-image retrieval performance of vision-language model should be evaluated in a multi-domain setting, characterized by a highly diverse pool of candidate images. Considering the limitations of zero-shot and finetuning methods in this context, we propose a novel Episodic Few-Shot Adaptation (EFSA) method, designed to enhance robustness against hard negatives in open-domain text-to-image retrieval tasks. By leveraging the top-$k$ candidate images along with synthetic captions generated for them, EFSA dynamically adapts to both domain- and sample-specific features, used to re-rank the top candidates and bring the ground-truth image to the very first rank. This approach consistently outperforms traditional fine-tuning and zero-shot retrieval methods across various benchmarks, demonstrating its effectiveness in mitigating domain-specific challenges and distributional shifts.

### v3

@@ 6 Conclusion
In this paper, we argue that the text-to-image retrieval performance of vision-language model should be evaluated in a multi-domain setting, characterized by a highly diverse pool of candidate images. Considering the limitations of zero-shot and finetuning methods in this context, we propose a novel Episodic Few-Shot Adaptation (EFSA) method, designed to enhance robustness against hard negatives in open-domain text-to-image retrieval tasks. By leveraging the top-$k$ candidate images along with synthetic captions generated for them, EFSA dynamically adapts to both domain- and sample-specific features, used to re-rank the top candidates and bring the ground-truth image to the very first rank. This approach consistently outperforms traditional fine-tuning and strong baselines across various benchmarks, demonstrating its effectiveness in mitigating domain-specific challenges and distributional shifts.
@@ 7 Experiments with SigLIP
8 Effect of Caption Generation Prompts
Figure 6 shows how retrieval performance changes with different prompts for generating image captions. We tested prompts that varied in length constraints, from no word limit to a maximum of 10, 20, 30, or 40 words. Overall, the choice of prompt has less than 1 point impact across Flickr30k and ArtCap. Performance improves when captions increase from 10 to 20 words but starts to decline as the word count goes beyond 20.
Figure 6: Effects of various caption generation prompts.
@@ 9 Qualitative Analysis
10 Storage Overhead Analysis
We compare the storage requirements introduced by EFSA—due to storing additional top-$k$ captions—with the zero-shot baseline that stores only image embeddings.
Baseline Storage:
We estimate the storage requirement of the zero-shot T2I baseline as:
Total Storage $=\langle\textit{size of retrieval pool}\rangle$
$\quad\times\langle\textit{bytes per cached image rep}\rangle.$ (3)
Each cached image embedding is a 768-dimensional float32 vector (as used in CLIP-base), requiring:
$768\times 4=3072\ \text{bytes per image}.$ (4)
EFSA Caption Overhead:
EFSA adds one generated caption per image, stored as a sequence of CLIP vocabulary token IDs, each encoded as a uint16. Assuming an average of 30 tokens per caption, the additional storage required is:
[…]
Relative Overhead:
The relative increase in storage per image is given by:
$\frac{60}{3072}\approx 0.0195\ (\approx 2\%).$ (6)
Thus, EFSA introduces a minimal storage overhead of approximately 2% compared to the zero-shot baseline, while providing measurable gains in both open- and closed-domain retrieval performance.
@@ 10 Storage Overhead Analysis
11 Effect of Loss Function
Loss Function COCO ArtCap
R@1 R@5 R@10 R@1 R@5 R@10
Hinge 30.15 50.78 57.79 10.91 22.73 27.47
Contrastive 28.23 49.20 56.47 10.33 21.66 26.61
Combined 30.14 50.96 57.82 11.13 22.83 27.50
12 Impact of False Negatives
In an open-domain setting, a given query $x$ may have multiple valid image matches $Y$, even though our evaluation setup only designates a single ground-truth. For instance, queries from COCO may have equally valid counterparts in closely related datasets such as Flickr, which are then characterized as false negatives. This ambiguity makes evaluation nuanced. Nevertheless, EFSA consistently improves Recall@1 with respect to the designated ground-truth targets, suggesting that the false negative issue, while possible, is not ubiquitous. In cases where it does interfere with the interpretation of Recall@1, we can additionally rely on Recall@5 and Recall@10, which reflect whether the ground-truth is still being successfully promoted to the top ranks despite the presence of potential false negatives.